# AI Ethics Principles in Practice: Perspectives of Designers and Developers

### v7

@@ 3. $\bullet$
Privacy Protection and Security
AI systems should respect and uphold privacy rights and data protection, and ensure the security of data.
Reliability and Safety
AI systems should reliably operate in the context of their intended purpose throughout their lifecycle.
Transparency and Explainability
There should be transparency and responsible disclosure to ensure people know when they are being significantly impacted by an AI system, and can find out when an AI system is engaging with them. Explainability includes what the AI system is doing and why, and may include the system’s processes and input data.
Fairness
AI systems should be inclusive and accessible, and should not involve or result in unfair discrimination against individuals or groups.
Contestability
When an AI system significantly impacts a person, group or environment, there should be a timely process to allow people to challenge the use or output of the system.
Accountability
Those responsible for the various phases of the AI system lifecycle should be identifiable and accountable for the outcomes of the system, and human oversight of AI systems should be enabled.
Human-centred Values
AI systems should respect human rights, diversity, and the autonomy of individuals.
Human, Social and Environmental (HSE) Well-being
AI systems should benefit individuals, society, and the environment.
@@ IV Results
Privacy Protection and Security
17 / 21 (81%)
Reliability and Safety
19 / 21 (90%)
Transparency and Explainability
18 / 21 (86%)
Fairness
10 / 21 (48%)
Contestability
 8 / 21 (38%)
Accountability
13 / 21 (62%)
Human-Centred Values
 3 / 21 (14%)
HSE Well-being
11 / 21 (52%)
@@ VII Conclusion
A notable challenge that emerged was balancing the tensions and trade-offs that were inherent to using the principles [^1] In follow-up work we have examined various approaches for resolving the tensions and trade-offs [57].. Tensions were highlighted between the practical approaches used for implementing privacy and security, transparency and explainability, as well as accuracy. For example, techniques may be used to develop an AI system that provides highly accurate outputs with the caveat that those outputs are not explicitly explainable. Similarly, the development of a highly explainable AI system may be vulnerable to privacy risks. In these cases, a choice can be made to prioritise one set of values over another [72].

### v8

@@ 3. $\bullet$
Privacy Protection and Security AI systems should respect and uphold privacy rights and data protection, and ensure the security of data.
Reliability and Safety AI systems should reliably operate in the context of their intended purpose throughout their lifecycle.
Transparency and Explainability There should be transparency and responsible disclosure to ensure people know when they are being significantly impacted by an AI system, and can find out when an AI system is engaging with them. Explainability includes what the AI system is doing and why, and may include the system’s processes and input data.
Fairness AI systems should be inclusive and accessible, and should not involve or result in unfair discrimination against individuals or groups.
Contestability When an AI system significantly impacts a person, group or environment, there should be a timely process to allow people to challenge the use or output of the system.
Accountability Those responsible for the various phases of the AI system lifecycle should be identifiable and accountable for the outcomes of the system, and human oversight of AI systems should be enabled.
Human-centred Values AI systems should respect human rights, diversity, and the autonomy of individuals.
Human, Social and Environmental (HSE) Well-being AI systems should benefit individuals, society, and the environment.
@@ IV Results
Privacy Protection and Security 17 / 21 (81%)
Reliability and Safety 19 / 21 (90%)
Transparency and Explainability 18 / 21 (86%)
Fairness 10 / 21 (48%)
Contestability  8 / 21 (38%)
Accountability 13 / 21 (62%)
Human-Centred Values  3 / 21 (14%)
HSE Well-being 11 / 21 (52%)
@@ VII Conclusion
A notable challenge that emerged was balancing the tensions and trade-offs that were inherent to using the principles111 In follow-up work we have examined in more detail various trade-offs and frameworks for resolving them [57, 58]. . Tensions were highlighted between the practical approaches used for implementing privacy and security, transparency and explainability, as well as accuracy. For example, techniques may be used to develop an AI system that provides highly accurate outputs with the caveat that those outputs are not explicitly explainable. Similarly, the development of a highly explainable AI system may be vulnerable to privacy risks. In these cases, a choice can be made to prioritise one set of values over another [73].